CH4 + 2O2 = CO2 + 2H2O

How many grams of O2 is needed to produce 37.5 g of CO2.



1
Expert's answer
2022-04-25T05:10:03-0400

CH4 + 2O2 = CO2 + 2H2O

m(O2) = (m(CO2.)*2Mr(O2))/Mr(CO2.)


m(O2) = (37.5 g * (2*32 g/moll)) / 44 g/moll = 54.5 g